1. A particle is travelling along a straight path described as = 0.75x . If its position along the x axis is
x= (8t)m, where t is in seconds, determine its speed when t=2s.
1
Expert's answer
2020-08-26T11:23:22-0400

Particle A particle is travelling along a straight path described as "y= 0.75x" .

And, "x=8t \implies y=0.75\times 8t=6t"

So particle path is given by "s=(8t) \ \hat{i} +(6t)\ \hat{j}"

Velocity of Particle "(v)=\frac{ds}{dt}=\frac{d}{dt} (8t) \hat{i} +\frac{d}{dt} (6t) \hat{j}=8\hat{i}+ 6\hat{j}\ m/s"

Or Magnitude will be "|v| =\sqrt{8^2+6^2}=10 \ m/s"